# Forecasting Events in Soccer Matches Through Language

### v1

@@ 3.5. Implementation
4. Experiments
4.1. Training the LEMs
@@
  • K=1: a model which only uses the the previous event as the game state, similar to the original proposal of LEM.
  • K=1s: a lite model trained with a significantly smaller number of parameters (6̃x less parameters than K=1).
@@
The performance of the three models is presented in Table 3.
Table 3. Comparative performance metrics of baseline and LEM predictive models across various variables and metrics. The table showcases accuracy (ACC) and F1-scores for categorical outcomes (Event Type, Goal, Accurate, Home), as well as Mean Absolute Error (MAE) and R-squared (R2) values for continuous variables (Time Elapsed, X, Y). This comprehensive analysis highlights the improvements in predictive accuracy and precision when utilizing different configurations of the LEM model compared to the baseline (BL). The LEM column corresponds to the original proposal (Mendes-Neves et al., 2024).
@@
Our proposed LEMs significantly improve soccer event prediction accuracy compared to a traditional baseline model and the previous iteration of LEMs. Table 3 reveals substantial gains across various metrics, including event type, accuracy prediction, and spatial coordinates. Only in two variables, isHome and TimeElapsed, our proposal did not improve over the previous iteration. Moreover, some improvements are substantial: (1) event type accuracy increased 6.5% for the K=3 model, (2) spatial coordinates X and Y decreased error by 24% and 28% respectively.
The results regarding the Goal/Accurate variables offer a glimpse into the trade-off this approach has against the original LEM proposal. The greater granularity of predictions, which are now predicted individually rather than massively, leads to increased performance in the variable that is predicted later, at a performance cost of the variable that was previously predicted. The original LEM proposal leads to better predictive performance. However, the methodology requires multiple variables to be forecasted simultaneously. This is where our current proposal allows for a significant leap. We can now forecast all variables sequentially, increasing the amount of information available in each event prediction, making the predicted events more cohesive than in the original approach.
We also ran tests to evaluate the inference speed of each algorithm. Comparing to K=1, K=1s offers a 21% increase in inference speed, while K=3 slows down inference by 62%.
4.2. Model Inspection
@@
Figure 3. The probability of transitioning from current location x,y to the next location x,y. The pattern contains two behaviors: (1) the positive correlation between the current coordinates and the next coordinate, as the next event performed by the same team is expected to be close to the current event, and (2) a negative correlation caused by when the next event is performed by the opposite team, as the coordinate axis changes to the opposition’s perspective.
@@
4.3. Computing Situational Expected Goals Maps
The expected goals metric (xG) is a widely used measure of opportunity quality in soccer. It measures the probability of a given shot situation leading to a goal. Figure 4 shows the expected goals map from two different game states. The first game state happens after a pass in the coordinates (80, 50), while the second happens after a cross in the coordinates (90, 80). For the K=3 model, both actions are preceded by two passes from coordinates (20, 50) and (50, 50), emulating a fast transition situation.
@@ 4.3. Computing Situational Expected Goals Maps
Figure 4. The situational expected goals maps calculated across the different models. For each case we simulated 1.000.000 shots for each input. Then, we calculate the percentage of shots leading to a goal for each location, which is used to plot the figures.
Several insights are visible in 4. First, we observe that there are data points with high distance to the goal and anomalous values for expected goals. These anomalies occur due to a bias in event data: when a cross goes in the direction of the goal, it is labeled as a shot if it ends in a goal but is still labeled as a cross if it gets claimed by the keeper. This discrepancy leads to divergences between real goal expectations and our model’s goal expectation.
@@
Interesting differences arise between K=1 and K=3. As both models are large enough to learn the patterns in data, the larger context in the K=3 model learns different patterns from the K=1 model. The K=3 model has access to two previous actions that indicate progressive passes from the shooting team. This means that the model is expecting a shot from a situation that is likely to be a fast transition. This added information increases the precision of the maps from the K=1 model that is unaware of this context.
4.4. Forecasting Short-term Probabilities
@@
LEMs also have the ability to forecast probabilities at multiple time scales. In soccer, we define short-term probability as the probability of scoring a goal in a short period after the event.
@@
The most famous use case for short-term probabilities are the momentum indicators in live score apps (e.g., Sofascore). Figure 5 shows an example of a momentum indicator.
@@
Figure 5. The visualization of a match momentum indicator built using the K=1 model. The data corresponds to the Real Madrid - Barcelona, December 23, 2017.
4.5. Forecasting Long-term Probabilities
@@
In opposition to short-term probabilities, long-term probabilities refers to a large time scale, usually to predict the outcome of a full match. With the simulations ran on LEMs, we can calculate the frequency of outcomes across a range of variables - everything from who wins (see Figure 6, to goals (see Figure 7, corners and others.
@@
Figure 6. The in-game probabilities calculated using LEMs for the game Real Madrid - Barcelona, December 23, 2017. The second half starts with a balance in probabilities and shifts abruptly every time Barcelona scores a goal. It is also visible in the image the short-term fluctuations provoked by events in the match.
@@
Figure 7. The in-game probabilities for whether the game will have under or over 2.5 goals, calculated using LEMs for the game Real Madrid - Barcelona, December 23, 2017. The probability of the current winning bet (under) increases as time passes, but sharply decreases when goals are scored. After the third goal, the outcome of this market is certain and the probabilities sharply move to 0 and 1 respectively.
Figure 7. The in-game probabilities for whether the game will have under or over 2.5 goals, calculated using LEMs for the game Real Madrid - Barcelona, December 23, 2017. The probability of the current winning bet (under) increases as time passes, but sharply decreases when goals are scored. After the third goal, the outcome of this market is certain and the probabilities sharply move to 0 and 1 respectively.
@@
The forecasted probabilities for both game outcome and the probability of the game ending with over or under 2.5 goals follow the expected pattern. The current outcome of the match increases its probability until a score change happens. When a goal happens, the probabilities sharply move, benefiting the team that scored.
@@
Both Figures 6 and 7 look like they have a lot of noise on them. The noise represents the small fluctuations of probabilities that happen due to the flow of the match: as a team becomes likelier to score, their probabilities increase at a much smaller scale than the increase they have when scoring a goal. Therefore, the probability noise actually resembles the changes in probability between events.
@@
One of the advantages of using LEMs for such purposes is that with the same simulations we can calculate an wide range of probabilities. For each new variable we want to forecast, we do not need to run a new simulation. We can reuse the generated event data from previous simulations and extract the probabilities. As for the limitations, the model cannot capture the shift in probabilities due to a red card in the 63rd minute, right before Barcelona’s second goal. This factor is absent from the context given to LEMs.
4.6. VAEP
@@
When estimating both short-term and long-term probabilities in soccer matches, we enable the calculation of action values via the Valuing Actions by Estimating Probabilities (VAEP) framework (Decroos et al., 2019). VAEP defines the value of an action as the resultant increase in the probability of scoring minus the increase in the probability of conceding. Using this framework, we calculated the VAEP value for the actions before the last goal in the Real Madrid - Barcelona, December 23, 2017. These results are presented in Table 4. We compare VAEP scores with ST/10 values, which assess the short-term scoring impact of actions within the time frame of 10 actions, to understand their alignment. Furthermore, we examine long-term probabilities (LT/inf) and a modified version (LT*/inf) calculated with a hypothetical 0-0 score to evaluate their effectiveness in different contexts. All values use the K=1 model.
@@
Table 4. The table compares the VAEP measures of 4 different approaches. The VAEP column represents the original VAEP approach (Decroos et al., 2019). ST/10 stands for short-term probability over a horizon of 10 events, which are parameters that allow a fair comparison with VAEP. LT/inf stands for long-term probability over an infinite horizon - this means that the game in simulated until the ¡GAME_OVER¿ event type appears in the simulation. LT*/inf modifies the current score to 0-0, which substantially changes the context of the measures. Both LT approaches weight outcomes by the amount of points it yields to the team (3 for a win, 1 for a draw, 0 for a loss) since soccer matches do not have binary outcomes.
@@
One action where VAEP and ST/10 disagree is the value of the dribble. The difference is explained by the better context given in the original VAEP probability estimator - the estimator knows where the dribble ends, while LEMs currently do not accept this information as input. Therefore, the added information of VAEP leads to a different valuation for the dribble. The initial LT/inf values, considering the entire game and its outcome, underestimate the value of the final shot since the match was already decided. However, when we adjust for a hypothetical 0-0 score (LT*(inf), the value of the shot aligns more accurately with the expected VAEP impact, although using a different scale.
These results are very important from the point of view of player valuation. Currently, the most relevant metrics for player valuation are aggregating the values of their events. Therefore, by evaluating events we are effectively evaluating the players.
5. Conclusion
@@
Future research focusing on contextual enrichment has the potential to unlock even greater accuracy and broader applicability of LEMs. Exploring advanced architectures like RNNs and Transformers can also propel this approach to new heights. This approach is a significant step forward in developing LEMs, opening doors for groundbreaking applications in soccer research. The possibilities are vast, from enhanced tactical analysis to betting applications.
Acknowledgements.

### v2

@@ 3.5 Implementation and Reproducibility
4 Experiments
4.1 Training the LEMs
@@
  • K=1: a model that only uses the previous event as the game state, similar to the original proposal of LEM.
  • K=1s: a lite model trained with a significantly smaller number of parameters (around 6x fewer parameters than K=1).
@@
The performance of the three models is presented in Table 3. We also report the results of a baseline (BL), which uses the majority class as a prediction in classification and uses the average values of the variable in regression.
Table 3: Comparative performance metrics of baseline and LEM predictive models across various variables and metrics. The table showcases accuracy (ACC) and F1-scores for categorical outcomes (Event Type, Goal, Accurate, Home), as well as Mean Absolute Error (MAE) and R-squared (R2) values for continuous variables (Time Elapsed, X, Y). This comprehensive analysis highlights the improvements in predictive accuracy and precision when utilizing different configurations of the LEM model compared to the baseline (BL). The LEM column corresponds to the original proposal [11].
@@
Our proposed LEMs significantly improve soccer event prediction accuracy compared to a traditional baseline model and the previous iteration of LEMs. Table 3 reveals substantial gains across various metrics, including event type, accuracy prediction, and spatial coordinates. Only in two variables, isHome and TimeElapsed, our proposal did not improve over the previous iteration. Moreover, some improvements are substantial: (1) event type accuracy increased by 6.5% for the K=3 model, and (2) spatial coordinates X and Y decreased error by 24% and 28%, respectively.
The results regarding the Goal/Accurate variables offer a glimpse into this approach’s trade-off against the original LEM proposal. The greater granularity of predictions, which are now predicted individually rather than massively, leads to increased performance in the variable predicted later, at a performance cost of the previously predicted variable. The original LEM proposal leads to better predictive performance. However, the methodology requires multiple variables to be forecasted simultaneously. This is where our current proposal allows for a significant leap. We can now forecast all variables sequentially, increasing the information available in each event prediction and making the predicted events more cohesive than the original approach.
We also ran tests to evaluate the inference speed of each algorithm. Compared to K=1, K=1s offers a 21% increase in inference speed, while K=3 slows down inference by 62%.
4.2 Model Inspection
@@
Figure 2: The probability of transitioning from current location x,y to the next location x,y. The pattern contains two behaviors: (1) the positive correlation between the current coordinates and the next coordinate, as the next event performed by the same team is expected to be close to the current event, and (2) a negative correlation caused by when the next event is performed by the opposite team, as the coordinate axis changes to the opposition’s perspective.
@@
4.3 Computing Situational Expected Goals Maps
The expected goals metric (xG) is a widely used measure of opportunity quality in soccer. It measures the probability of a given shot situation leading to a goal. Figure 3 shows the expected goals map from two different game states. The first game state happens after a pass in the coordinates (80, 50), while the second occurs after a cross in the coordinates (90, 80). For the K=3 model, both actions are preceded by two passes from coordinates (20, 50) and (50, 50), emulating a fast transition situation.
@@ 4.3 Computing Situational Expected Goals Maps
Figure 3: The situational expected goals maps calculated across the different models. For each case, we simulated 1.000.000 shots for each input. Then, we calculate the percentage of shots leading to a goal for each location, which is used to plot the figures.
Several insights are visible in 3. First, we observe data points with a high distance from the goal and abnormal values for expected goals. These anomalies occur due to a bias in event data: when a cross goes toward the goal, it is labeled as a shot if it ends in a goal but is still labeled as a cross if it gets claimed by the keeper. This discrepancy diverges between real goal expectations and our model’s goal expectations. Nonetheless, this data bias is well reported in the literature and recurring across different datasets [9].
@@
Interesting differences arise between K=1 and K=3. As both models are large enough to learn the patterns in data, the larger context in the K=3 model learns different patterns from the K=1 model. The K=3 model has access to two previous actions that indicate progressive passes from the shooting team. This information helps the model expect a shot from a situation likely to be a fast transition. The added information increases the precision of the maps from the K=1 model that is unaware of this context.
4.4 Forecasting Short-term Probabilities
@@
LEMs also can forecast probabilities at multiple time scales. In soccer, we define short-term probability as the probability of scoring a goal shortly after the event.
@@
The most famous use case for short-term probabilities is the momentum indicators in live score apps (e.g., Sofascore). Figure 4 shows an example of a momentum indicator.
@@
Figure 4: The visualization of a match momentum indicator built using the K=1 model. The data corresponds to the Real Madrid - Barcelona, December 23, 2017.
4.5 Forecasting Long-term Probabilities
@@
In contrast to short-term probabilities, long-term probabilities refer to a large time scale, usually to predict the outcome of a full match. With the simulations ran on LEMs, we can calculate the frequency of outcomes across a range of variables - everything from who wins (see Figure 5, to goals (see Figure 6, corners and others.
@@
Figure 5: The in-game probabilities calculated using LEMs for the game Real Madrid - Barcelona, December 23, 2017. The second half starts with a balance in probabilities and shifts abruptly every time Barcelona scores a goal. The short-term fluctuations provoked by events in the match are also visible in the image.
@@
Figure 6: The in-game probabilities for whether the game will have under or over 2.5 goals, calculated using LEMs for Real Madrid - Barcelona, December 23, 2017. The probability of the current winning bet (under) increases as time passes but sharply decreases when goals are scored. After the third goal, the outcome of this market is certain, and the probabilities sharply move to 0 and 1, respectively.
@@
The forecasted probabilities for both game outcome and the probability of the game ending with over or under 2.5 goals follow the expected pattern. The current result of the match increases its probability until a score change happens. When a goal occurs, the probabilities sharply move, benefiting the team that scored.
@@
Figures 5 and 6 look like they have a lot of noise. The noise represents the slight fluctuations of probabilities due to the match flow: as a team becomes likelier to score, their probabilities increase at a much smaller scale than their increase when scoring a goal. Therefore, the probability noise resembles the changes in probability between events.
@@
One of the advantages of using LEMs for such purposes is that with the same simulations, we can calculate a wide range of probabilities. We do not need to run a new simulation for each new variable we want to forecast. We can reuse the generated event data from previous simulations and extract the probabilities. As for the limitations, the model cannot capture the shift in probabilities due to a red card in the 63rd minute, right before Barcelona’s second goal. This factor is absent from the context given to LEMs.
4.6 VAEP
@@
When estimating both short-term and long-term probabilities in soccer matches, we enable calculating action values via the Valuing Actions by Estimating Probabilities (VAEP) framework [1]. VAEP defines the value of an action as the resultant increase in the probability of scoring minus the increase in the probability of conceding. Using this framework, we calculated the VAEP value for the actions before the last goal in the Real Madrid - Barcelona, December 23, 2017. These results are presented in Table 4. We compare VAEP scores with ST/10 values, which assess the short-term scoring impact of actions within the time frame of 10 actions, to understand their alignment. Furthermore, we examine long-term probabilities (LT/inf) and a modified version (LT*/inf) calculated with a hypothetical 0-0 score to evaluate their effectiveness in different contexts. All values use the K=1 model.
@@
Table 4: The table compares the VAEP measures of 4 different approaches. The VAEP column represents the original VAEP approach [1]. ST/10 stands for short-term probability over a horizon of 10 events, which are parameters that allow a fair comparison with VAEP. LT/inf stands for long-term probability over an infinite horizon - this means that the game in simulated until the <GAME_OVER> event type appears in the simulation. LT*/inf modifies the current score to 0-0, substantially changing the measures’ context. Both LT approaches weight outcomes by the amount of points it yields to the team (3 for a win, 1 for a draw, 0 for a loss) since soccer matches do not have binary outcomes.
@@
One action where VAEP and ST/10 disagree is the value of the dribble. The difference is explained by the better context in the original VAEP probability estimator - the estimator knows where the dribble ends, while LEMs currently do not accept this information as input. Therefore, the added information of VAEP leads to a different valuation for the dribble. The initial LT/inf values, considering the entire game and its outcome, underestimate the value of the final shot since the match was already decided. However, when we adjust for a hypothetical 0-0 score (LT*(inf), the value of the shot aligns more accurately with the expected VAEP impact, although using a different scale.
These results are very important from the point of view of player valuation. Currently, the most relevant metrics for player valuation are aggregating the values of their events. Therefore, by evaluating events, we are effectively assessing the players.
5 Conclusion
@@
Future research focusing on contextual enrichment has the potential to unlock even greater accuracy and broader applicability of LEMs. Exploring advanced architectures like RNNs and Transformers can also propel this approach to new heights. This approach is a significant step forward in developing LEMs, opening doors for groundbreaking applications in soccer research. The possibilities are vast, from enhanced tactical analysis to match prediction applications.
{ack}